OpenStudy (anonymous):

Rewrite the following equation in slope- intercept form. x+y=5 A. Y=x-5 B. Y= x+ 5 C. Y=-x+5 D. Y=-x-5

Nnesha (nnesha):

y=mx+b is slope intercept form where m is slope and b is y-intercept so solve x+y=5 for y

OpenStudy (anonymous):

So, B?

Nnesha (nnesha):

move `x` to the right side how would you cancel out x from left side ?

OpenStudy (anonymous):

Subtract it from both sides

Nnesha (nnesha):

yes right subtract x form side so what would be the answer?

OpenStudy (anonymous):

Y=-x+5?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Thanks!!